# The Generalization Ridge: Information Flow in Natural Language Generation

### v2

@@ Abstract
Transformer-based language models have achieved state-of-the-art performance in natural language generation (NLG), yet their internal mechanisms for synthesizing task-relevant information remain insufficiently understood. While prior studies suggest that intermediate layers often yield more generalizable representations than final layers, how this generalization ability emerges and propagates across layers during training remains unclear. To address this gap, we propose InfoRidge, an information-theoretic framework, to characterize how predictive information—the mutual information between hidden representations and target outputs—varies across depth. Estimating this quantity enables us to trace the flow of task-relevant information throughout the model during training. Our experiments across various models and datasets reveal a consistent non-monotonic trend: predictive information peaks in upper-middle layers—forming a generalization ridge—before declining in final layers, reflecting a transition between generalization and memorization. To further investigate this phenomenon, we introduce residual scaling coefficients—trainable scalar parameters applied to each residual block—which serve as functional probes for assessing the relative importance of individual transformer layers. These coefficients reveal that, under distribution shift, models downweight final layers and increasingly rely on intermediate layers, highlighting their role in generalization. Together, these findings offer new insights into the internal mechanisms of transformers and underscore the critical role of intermediate layers in supporting generalization.
@@ 1 Introduction
A growing body of research has shown that intermediate layers in deep neural networks often surpass final layers in terms of representational quality and generalization performance (Liu et al., 2019b; Voita et al., 2019; Ansuini et al., 2019; Ahrens et al., 2023; Uselis & Oh, 2025). In language models, intermediate layers often encode richer semantic and more robust features than final layers (Fan et al., 2024; Jin et al., 2024; Skean et al., 2025). However, questions still remain: In NLG, how does information evolve across layers during training, and how are different layers of the network functionally organized to support generalization versus memorization?
@@
To further validate this interpretation, we introduce residual scaling coefficients—learnable scalar parameters $\beta_{\ell}$ applied to each residual block—while keeping all other model weights frozen, drawing inspiration from prior work on layer-wise adaptation (Liu et al., 2019a; Menghani et al., 2024). A higher $\beta_{\ell}$ value indicates greater reliance on the corresponding layer’s output during prediction. These coefficients act as functional probes, revealing how the model redistributes layer-wise importance under different data distribution. Under in-distribution training, deeper layers retain higher residual weights, reflecting the model’s reliance on memorized, task-specific features. When evaluated under distribution shift, models have lower reliance on the final layer and shifts importance toward the intermediate layers.
To understand the formation of the generalization ridge, we analyze both attention patterns and model capacity. Attention analysis shows ridge layers attend to tokens that capture broadly useful features, aligning with the information peak. Beyond attention patterns, we also investigate the conditions under which the ridge emerges. Our depth ablation results show that the ridge only emerges beyond a certain depth threshold. Below this threshold, predictive information increases monotonically, indicating that sufficient capacity is a prerequisite for generalization ridge to emerge.
@@
3. We introduce residual scaling coefficients as trainable indicators of how models shift representational focus across layers during training, providing a causal, adaptive measure of model behavior under distribution shifts.
@@ 2 Related Work
Understanding how information is encoded and transformed across layers has been studied through probing classifiers (Alain & Bengio, 2016), attention flow (Vig & Belinkov, 2019), and information-theoretic approaches such as the information bottleneck (Shwartz-Ziv & Tishby, 2017), mutual information estimation (Goldfeld, 2019), and matrix-based entropy (Giraldo et al., 2014), offering different lenses to quantify representational capacity, abstraction, and invariance across layers.
@@
A growing body of research has shown that intermediate layers in deep networks often outperform final layers in terms of representational quality and task performance (Ansuini et al., 2019; Yosinski et al., 2014; Uselis & Oh, 2025; Ahrens et al., 2023; Ando et al., 2023). In language models, mid-depth layers tend to capture richer semantic or robust features than output layers (Liu et al., 2019b; Voita et al., 2019; Jin et al., 2024; Fan et al., 2024). Transformer representations have long been observed to follow a structured progression from syntactic to semantic information, as shown by classical probing studies on linguistic knowledge and the reconstruction of the NLP pipeline (Liu et al., 2019b; Tenney et al., 2019). These findings challenge the assumption that deeper layers always yield better representations.
@@
This pattern holds across settings such as transfer learning (Yosinski et al., 2014), continual learning (Ahrens et al., 2023), and out-of-distribution generalization (Uselis & Oh, 2025). Furthermore, recent work has evaluated representation quality using entropy, curvature, and invariance (Skean et al., 2025), while other studies have examined embedding drift and representational geometry (Merchant et al., 2020; Dar et al., 2022), analyzed memorization and factual recall (Haviv et al., 2022; Yu et al., 2023), and introduced causal perspectives on layer importance through mediation analysis and targeted interventions (Vig et al., 2020; Meng et al., 2022). Training-dynamics studies investigated how earlier models develop and refine semantic features across depth, providing an additional perspective on the evolution of layer-wise representations (Merchant et al., 2020; Kumar et al., 2023). Correlational probes (e.g., linear probes (Alain & Bengio, 2016)) measure only whether a feature can be decoded from a representation, which reflects correlation but not causal influence. In contrast, causal methods intervene on internal activations to test how changing a component alters the model’s prediction, thereby identifying true causal contribution rather than mere feature presence. Our residual-scaling approach aligns with this causal perspective at the layer level.
@@
However, the underlying causes and functional role of this phenomenon remain only partially understood, motivating further investigation. Our work addresses this gap by tracing the evolution of predictive information throughout training and establishing a clear connection between predictive information flow and generalization. We reveal a consistent non-monotonic peak in the middle layers—termed the generalization ridge—that reflects a meaningful transition in representational focus and aligns with stronger generalization behavior. Additionally, unlike prior work focused on classification tasks, we extend the analysis of information flow to generation tasks by quantifying the mutual information between hidden states and the next token. This enables us to understand how generalization and memorization dynamics evolve for next-token generation setting during training from an information-theoretic perspective.
@@ 3 InfoRidge: Information Estimation Framework
$Z_{\ell}=\{z^{(\ell)}_{i}\}_{i=1}^{N},\quad Y=\{y_{i}\}_{i=1}^{N},\quad\Delta Z_{\ell}=\{\delta z_{i}^{(\ell)}\}_{i=1}^{N},$
@@
We apply the matrix-based framework (Giraldo et al., 2014) to estimate mutual information. Let $\mathcal{U}$ be a random variable, from which we draw a set of vectors $U=\{\mathbf{u}_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d}$. A positive-definite Gram matrix $G_{U}\in\mathbb{R}^{N\times N}$ is computed using a Gaussian kernel $\kappa$ with bandwidth set to 1 and the matrix is then trace-normalized to satisfy $\operatorname{tr}(G_{U})=1$. The matrix-based Rényi entropy (with order $\alpha=1$) is then given by:
@@
$I(\mathcal{U};\mathcal{V})\approx I(G_{\mathcal{U}};G_{\mathcal{V}})=H(G_{\mathcal{U}})+H(G_{\mathcal{V}})-H(G_{\mathcal{U}}\circ G_{\mathcal{V}}),$ (2)
where $\circ$” denotes the Hadamard (elementwise) product. Mathematical details are in Appendix A.
@@ 4 Experimental Setup
We assess model behavior across three tasks casted into NLG problems: CLUTRR (Sinha et al., 2019), a relational reasoning benchmark; ECQA (Aggarwal et al., 2021), a commonsense QA benchmark; and Synthetic Arithmetic, a controlled dataset designed to disentangle task-relevant signal from noise. Dataset and implementation details are in Appendix B and C.
@@ 5.1 Predictive Information: Information Peaks at Intermediate Layers
This discussion aims to provide an intuitive illustration of our generalization ridge hypothesis, highlighting how attention to task-relevant signal tokens shifts across layers (the generalizable information). Specifically, we computed average signal attention across layers, identified the layer with peak signal attention, and compared it to final-layer signal and last-token attention, results are in Table 2. Signal tokens are defined as task-relevant tokens that the model must focus on to solve the task—for the synthetic dataset, these are tokens that appear after the character ‘S’; for CLUTRR, kinship-related entities; and for ECQA, the token corresponding to the correct answer option. This allows us to quantify where in the network attention to semantically important information is concentrated. Our findings show a pattern that supports our hypothesis: (1) Mid-to-late layers peak in signal attention, coinciding with the predictive information ridge (Figure 2), indicating where generalizable representations are strongest. (2) Final layers show reduced attention to signal tokens, suggesting a shift toward memorization to specific data point rather than predictive information abstraction. For example, in Qwen-2.5-0.5B on ECQA, signal attention peaks at Layer 17 (0.2458) but drops to 0.0066 in the final layer, where last-token attention dominates (0.1697). Additional results are in Appendix F.
@@ 5.2 Incremental Information Gain: Information Concentrates at Intermediate Layers
To understand how information accumulates across the network, we compute Incremental Information Gain ($I(\Delta Z_{\ell};Y)$)—the mutual information between each residual transition and the target label embedding. As shown in Figure 5, the resulting layer-wise gains reveal that intermediate layers yield the highest information increases. This concentration of information gain further underscores their central role in encoding those task-relevant features that are essential for supporting generalization. For additional illustration, a detailed case study that decodes $\Delta z$ via the LM head is provided in Appendix G, highlighting fine-grained token-level shifts and variations observed across layers.
@@
The Incremental Information Gain analysis reveals a clear pattern: middle transformer blocks are key to encoding generalizable task-relevant information, thereby forming a generalization ridge. Conversely, later layers contribute little additional predictive signal, and in some cases, actively reduce alignment with the target embeddings. This diminishing contribution in later layers may suggest a shift from general reasoning to memorization of training-specific patterns. This trend underscores a fundamental trade-off in transformer training dynamics.
6 Residual Scaling Dynamics via Learnable $\beta$ Coefficients
@@
We introduce a residual scaling mechanism with learnable scalar coefficient parameters (Algorithm 1), inspired by prior work on adaptive residual modulation (Liu et al., 2019a; Menghani et al., 2024). Transformer architectures inherently employ residual connections to iteratively refine representations. To isolate and quantify the contribution of each transformer block, we scale these residual connections with layer-specific scaling factors $\beta_{\ell}\in\mathbb{R}_{\geq 0}$, with definition below. Each $\beta_{\ell}$ controls the strength of the residual contribution from layer $\ell$, enabling the model to adaptively emphasize or suppress specific blocks:
@@
Algorithm 1 Residual Scaling: Probing the Contribution of Transformer Blocks to the Generalization–Memorization Trade-off.
1:Pretrained Transformer with $L$ layers, dataset $\mathcal{D}$, learning rate $\eta$
2:Initialize $\beta_{1},\dots,\beta_{L}\leftarrow 1.0$ (trainable); freeze other weights
3:for each training step do
4:  Sample batch $(x,y)\sim\mathcal{D}$
5:  $z^{(0)}\leftarrow\text{Embedding}(x)$
6:  for $\ell=1$ to $L$ do
7:   $r^{(\ell)}\leftarrow\text{block}^{(\ell)}(z^{(\ell-1)})$
8:   $z^{(\ell)}\leftarrow z^{(\ell-1)}+\beta_{\ell}\cdot r^{(\ell)}$
9:  end for
10:  Compute loss $\mathcal{L}(z^{(L)},y)$
11:  Update $\beta_{1},\dots,\beta_{L}$ using gradient descent
12:end for
@@
Figure 6: Residual scaling coefficients $\beta_{\ell}$ across all transformer layers. ID training emphasizes later layers, while OOD training has lower final layer weight. Curve shows the mean across 5 independent run, and the shaded region denotes 1-$\sigma$ error bar.
@@
As shown in Figure 6, optimizing the residual scaling coefficient parameters on in-distribution data assigns relatively higher residual weights to the final transformer layers, indicating that in-distribution performance benefits from increased reliance on late-layer residual contributions. This observation suggests that these deeper layers capture refinements that are closely tied to the training distribution. In contrast, under out-of-distribution (OOD) settings, the learned coefficients systematically have relative lower contribution of the final layers compared to the ID settings, revealing that improved generalization is associated with downweighting late-stage residual signals. More results are in Appendix E. This pattern supports the interpretation that late-layer residuals become increasingly specialized in memorized patterns tied to the training distribution. Together, these findings provide empirical support for the generalization ridge hypothesis, demonstrating that information flow in transformers reflects a trade-off between generalizable signals and memorized, distribution-specific features.

### v3

@@ Abstract
Transformer-based language models have achieved state-of-the-art performance in natural language generation (NLG), yet their internal mechanisms for synthesizing task-relevant information remain insufficiently understood. While prior studies suggest that intermediate layers often yield more generalizable representations than final layers, how this generalization ability emerges and propagates across layers during training remains unclear. To address this gap, we propose InfoRidge, an information-theoretic framework, to characterize how predictive information—the mutual information between hidden representations and target outputs—varies across depth during training. Our experiments across various models and datasets reveal a consistent non-monotonic trend: predictive information peaks in intermediate layers—forming a generalization ridge—before declining in final layers, reflecting a transition between generalization and memorization. To further investigate this phenomenon, we conduct a set of complementary analyses that leverage residual scaling, attention pattern, and controlled model capacity to characterize layer-wise functional specialization. We further validate our findings with multiple-token generation experiments, verifying that the observed ridge phenomenon persists across decoding steps. Together, these findings offer new insights into the internal mechanisms of transformers and underscore the critical role of intermediate layers in supporting generalization.
@@ 1 Introduction
A growing body of research has shown that intermediate layers in deep neural networks often surpass final layers in terms of representational quality and generalization performance (Liu et al., 2019b; Voita et al., 2019; Ansuini et al., 2019; Ahrens et al., 2023; Uselis and Oh, 2025). In language models, intermediate layers often encode richer semantic and more robust features than final layers (Fan et al., 2024; Jin et al., 2024; Skean et al., 2025). However, questions still remain: In NLG, how does information evolve across layers during training, and how are different layers of the network functionally organized to support generalization versus memorization?
@@
We further extend InfoRidge to multi-token natural language generation. By treating the full output sequence as a structured target and measuring mutual information between hidden states and the long-form target distribution, we show that the generalization ridge persists across decoding steps. Despite increasing mutual information as generation progresses, the layer-wise structure remains stable: predictive information consistently peaks in early-to-intermediate layers and declines in later layers.
To understand the formation of the generalization ridge, we analyze attention patterns, model capacity and residual scaling coefficients. Attention analysis shows ridge layers attend to tokens that capture broadly useful features, aligning with the information peak. We also investigate the conditions under which the ridge emerges. Our depth ablation results show that the ridge only emerges beyond a certain depth threshold. Below this threshold, predictive information increases monotonically, indicating that sufficient capacity is a prerequisite for generalization ridge to emerge. We also introduce residual scaling coefficients $\beta_{\ell}$—learnable scalars applied to each residual block with all other weights frozen—as functional probes of layer-wise reliance (Liu et al., 2019a; Menghani et al., 2024). When evaluated under distribution shift, models have lower reliance on the final layer and shifts importance toward the intermediate layers.
@@
3. We analyze attention patterns, model capacity and residual scaling coefficients, providing causal evidences for the generalization ridge phenomenon.
@@ 2 Related Work
Understanding how information is encoded and transformed across layers has been studied through probing classifiers (Alain and Bengio, 2016), attention flow (Vig and Belinkov, 2019), and information-theoretic approaches such as the information bottleneck (Shwartz-Ziv and Tishby, 2017), mutual information estimation (Goldfeld, 2019), and matrix-based entropy (Giraldo et al., 2014), offering different lenses to quantify representational capacity, abstraction, and invariance across layers.
@@
A growing body of research has shown that intermediate layers in deep networks often outperform final layers in terms of representational quality and task performance (Ansuini et al., 2019; Yosinski et al., 2014; Uselis and Oh, 2025; Ahrens et al., 2023; Ando et al., 2023). In language models, mid-depth layers tend to capture richer semantic or robust features than output layers (Liu et al., 2019b; Voita et al., 2019; Jin et al., 2024; Fan et al., 2024). Transformer representations have long been observed to follow a structured progression from syntactic to semantic information, as shown by classical probing studies on linguistic knowledge and the reconstruction of the NLP pipeline (Liu et al., 2019b; Tenney et al., 2019). These findings challenge the assumption that deeper layers always yield better representations.
@@
This pattern holds across settings such as transfer learning (Yosinski et al., 2014), continual learning (Ahrens et al., 2023), and out-of-distribution generalization (Uselis and Oh, 2025). Furthermore, recent work has evaluated representation quality using entropy, curvature, and invariance (Skean et al., 2025), while other studies have examined embedding drift and representational geometry (Merchant et al., 2020; Dar et al., 2022), analyzed memorization and factual recall (Haviv et al., 2022; Yu et al., 2023), and introduced causal perspectives on layer importance through mediation analysis and targeted interventions (Vig et al., 2020; Meng et al., 2022). Training-dynamics studies investigated how earlier models develop and refine semantic features across depth, providing an additional perspective on the evolution of layer-wise representations (Merchant et al., 2020; Kumar et al., 2023). Correlational probes (e.g., linear probes (Alain and Bengio, 2016)) measure only whether a feature can be decoded from a representation, which reflects correlation but not causal influence. In contrast, causal methods intervene on internal activations to test how changing a component alters the model’s prediction, thereby identifying true causal contribution rather than mere feature presence. Our residual-scaling approach aligns with this causal perspective at layer level.
@@
However, the underlying causes and functional role of this phenomenon remain only partially understood, motivating further investigation. Our work addresses this gap by tracing the evolution of predictive information throughout training and establishing a clear connection between predictive information flow and generalization. We reveal a consistent non-monotonic peak in the middle layers—termed the generalization ridge—that reflects a meaningful transition in representational focus and aligns with stronger generalization behavior. Additionally, unlike prior work focused on classification tasks, we extend the analysis of information flow to generation tasks, enabling us to understand how generalization and memorization dynamics evolve for natural language generation setting from an information-theoretic perspective.
@@ 3 InfoRidge: Information Estimation Framework
$Z_{\ell}=\{z^{(\ell)}_{i}\}_{i=1}^{N}\!,\;Y=\{y_{i}\}_{i=1}^{N}\!,\;\Delta Z_{\ell}=\{\delta z_{i}^{(\ell)}\}_{i=1}^{N},$
@@
We apply the matrix-based framework (Giraldo et al., 2014) to estimate mutual information. Let $\mathcal{U}$ be a random variable, from which we draw a set of vectors $U=\{\mathbf{u}_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d}$. A positive-definite Gram matrix $G_{U}\in\mathbb{R}^{N\times N}$ is computed using a Gaussian kernel $\kappa$ with bandwidth set to 1 and the matrix is then trace-normalized to satisfy $\operatorname{tr}(G_{U})=1$. To assess robustness with respect to kernel selection, we additionally test the Laplacian and Polynomial kernels. While these kernels introduce quantitative shifts in the magnitude, the trend remains unchanged, see Appendix E. The matrix-based Rényi entropy (with order $\alpha=1$) is then given by:
@@
$I(\mathcal{U};\mathcal{V})$ $\approx I(G_{\mathcal{U}};G_{\mathcal{V}})$ (2)
$=H(G_{\mathcal{U}})+H(G_{\mathcal{V}})-H(G_{\mathcal{U}}\circ G_{\mathcal{V}}),$
where $\circ$” denotes the Hadamard product. Mathematical details are in Appendix A.
@@ 4 Experimental Setup
We assess model behavior across three tasks casted into NLG problems: CLUTRR (Sinha et al., 2019), a relational reasoning benchmark; ECQA (Aggarwal et al., 2021), a commonsense QA benchmark; CNN/DailyMail (See et al., 2017; Hermann et al., 2015), a summarization benchmark and Synthetic Arithmetic, a controlled dataset designed to disentangle task-relevant signal from noise. Dataset and implementation details are in Appendix B and C.
@@ 5.1 Predictive Information: Information Peaks at Intermediate Layers
This discussion aims to provide an intuitive illustration of our generalization ridge hypothesis, highlighting how attention to task-relevant signal tokens shifts across layers (the generalizable information). Specifically, we computed average signal attention across layers, identified the layer with peak signal attention, and compared it to final-layer signal and last-token attention, results are in Table 2. Signal tokens are defined as task-relevant tokens that the model must focus on to solve the task—for the synthetic dataset, these are tokens that appear after the character ‘S’; for CLUTRR, kinship-related entities; and for ECQA, the token corresponding to the correct answer option. This allows us to quantify where in the network attention to semantically important information is concentrated. Our findings show a pattern that supports our hypothesis: (1) Mid-to-late layers peak in signal attention, coinciding with the predictive information ridge (Figure 2), indicating where generalizable representations are strongest. (2) Final layers show reduced attention to signal tokens, suggesting a shift toward memorization to specific data point rather than predictive information abstraction. For example, in Qwen-2.5-0.5B on ECQA, signal attention peaks at Layer 17 (0.2458) but drops to 0.0066 in the final layer, where last-token attention dominates (0.1697).
@@ 5.2 Incremental Information Gain: Information Concentrates at Intermediate Layers
To understand how information accumulates across the network, we compute Incremental Information Gain ($I(\Delta Z_{\ell};Y)$)—the mutual information between each residual transition and the target label embedding. As shown in Figure 5, the resulting layer-wise gains reveal that intermediate layers yield the highest information increases. This concentration of information gain further underscores their central role in encoding those task-relevant features that are essential for supporting generalization. The Incremental Information Gain analysis reveals a clear pattern: middle transformer blocks are key to encoding generalizable task-relevant information, thereby forming a generalization ridge. Conversely, later layers contribute little additional predictive signal, and in some cases, actively reduce alignment with the target embeddings. This diminishing contribution in later layers may suggest a shift from general reasoning to memorization of training-specific patterns. This trend underscores a fundamental trade-off in transformer training dynamics. Additional results are in Appendix E.
@@
Residual Scaling Dynamics via Learnable $\beta$ Coefficients
@@
We introduce a residual scaling mechanism with learnable scalar coefficient parameters, inspired by prior work on adaptive residual modulation (Liu et al., 2019a; Menghani et al., 2024). Transformer architectures inherently employ residual connections to iteratively refine representations. To isolate and quantify the contribution of each transformer block, we scale these residual connections with layer-specific scaling factors $\beta_{\ell}\in\mathbb{R}_{\geq 0}$, with definition below. Each $\beta_{\ell}$ controls the strength of the residual contribution from layer $\ell$, enabling the model to adaptively emphasize or suppress specific blocks:
@@
Figure 6: Residual scaling coefficients $\beta_{\ell}$ across all layers. ID training emphasizes later layers, while OOD training has lower final layer weight. Curve shows the mean across 5 independent run, and the shaded region denotes 1-$\sigma$ error bar.
@@
As shown in Figure 6, optimizing the residual scaling coefficient parameters on in-distribution data assigns relatively higher residual weights to the final transformer layers, indicating that in-distribution performance benefits from increased reliance on late-layer residual contributions. This observation suggests that these deeper layers capture refinements that are closely tied to the training distribution. In contrast, under out-of-distribution (OOD) settings, the learned coefficients systematically have relative lower contribution of the final layers compared to the ID settings, revealing that improved generalization is associated with downweighting late-stage residual signals. Additional results are in Appendix E. This pattern supports the interpretation that late-layer residuals become increasingly specialized in memorized patterns tied to the training distribution. Together, these findings provide empirical support for the generalization ridge hypothesis, demonstrating that information flow in transformers reflects a trade-off between generalizable signals and memorized, distribution-specific features.
6 Extension to Multi-Token Senario
We further evaluate information flow under multi-token outputs using the CNN/DailyMail summarization dataset (1.0.0, test) (See et al., 2017; Hermann et al., 2015). We evaluate two pretrained base models: Qwen-2.5-0.5B and LLaMA-3.1-8B, generating 50 tokens per sample.
Kernel Extension for Variable-Length Sequences.
To handle multi-token targets, we extend our matrix-based mutual information (MI) estimation to variable-length sequences. Let $Y_{i}=\{y_{i}^{(t)}\}_{t=1}^{T_{i}}$ denote the target sequence for sample $i$. We employ the following kernel to compute sample-to-sample similarity:
$K(Y_{i},Y_{j})=\frac{1}{T_{i}T_{j}}\sum_{t=1}^{T_{i}}\sum_{t^{\prime}=1}^{T_{j}}\kappa\!\left(y_{i}^{(t)},y_{j}^{(t^{\prime})}\right),$
where $\kappa(\cdot,\cdot)$ is a Gaussian kernel. This formulation aggregates all pairwise token similarities between two target sequences into a single scalar, capturing the overall alignment between two token distributions.
We quantify how much information each layer carries about the gold answer by computing $I(Z_{\ell}^{(t)};Y)$ in Figure 7, where $Z_{\ell}$ denotes the hidden state at the last token position at decoding step $t$ and $Y$ represents the full target sequence.
(a) LLaMA on CNN/Daily for 50 generation steps.
(b) Qwen on CNN/Daily for 50 generation steps.
Figure 7: $I(Z_{\ell}^{(t)};Y)$ across layers (x-axis) and generation steps (left color bar), the right y-axis shows mutual information values. Values are plotted every ten steps. Lines denote the mean across five independent runs, with shaded regions indicating $\pm 1$ standard deviation.
Mutual information across generation steps.
We observe several consistent patterns in the mutual information profiles across generation steps. First, the overall magnitude of $I(Z_{\ell}^{(t)};Y)$ increases as generation progresses: later decoding steps tend to exhibit higher mutual information than earlier ones. This reflects the accumulation of information in the hidden representations as the model conditions on an increasingly informative prefix composed of its previously generated tokens. As generation progresses, the hidden state progressively approach the target sequence distribution.
Second, despite differences in magnitude, the layer-wise structure of the mutual information curves remains consistent across generation steps. In particular, $I(Z_{\ell}^{(t)};Y)$ peaks at early-to-intermediate layers and gradually declines toward later layers. The concentration of mutual information in early-to-intermediate layers reflects their role in forming task-relevant representations under autoregressive decoding. These layers integrate information from the input and the evolving context into relatively abstract features that are shared across decoding steps, such as semantic content and task-level structure. In contrast, representations in later layers are increasingly shaped by step-specific decoding requirements, including local token selection, memorized pretrained corpus-specific surface patterns and distribution sharpening. While these representations are critical for producing the next output token, later layers increasingly compress information into a lower-entropy, decision-focused subspace, leading to reduced mutual information with the target sequence.
Figure 8: Mutual information comparing stopwords/punctuation against content words.
Figure 8 compares $I(Z_{\ell};Y)$ when generating content words versus stopwords and punctuation. Both models exhibit similar layer-wise trends across the two token types, with curves largely overlapping throughout early and middle layers, suggesting that intermediate layers perform semantic processing regardless of output token type.
A divergence emerges in the final layers: content words show higher $I(Z_{\ell};Y)$ than stopwords. This reflects differences in lexical selection complexity. Content words draw from a large vocabulary space and require precise semantic discrimination, demanding more target information. Stopwords follow predictable syntactic patterns determined by local grammatical constraints, requiring less target information for selection.